How can I do the calculation
b. Concentration ofreactants c. Presence of a catalysts d. Concentration of products e. Surface area of solid reactant 4. Calculate the activation energy, Ea, for the reaction 2N205(g) → 4NO2(g) + O2(g) from the observed rate constants: k-3.46x10-5 s-1 at 25°C and k= 1.5x10-3 s-1 at 55°C e. 1.04 kJ/mol a. 1437 kJ/mol c. 0.101 kJ/mol b. 102 kJ/mol d. 9.62 kJ/mol CCOCOCCOCOCOOCLCOCCOO0008000
Answer #1

ln (K2/K1) = (Ea / R) (1/T1 - 1/T2)

T1 = 250C = 298 K K1 = 3.46 x 10-5 s-1

T2 = 550C = 328 K K2 = 1.5 x 10-3 s-1

ln (1.5 x 10-3 / 3.46 x 10-5) = (Ea / 8.314) (1/298 - 1/328)

3.77 = (Ea / 8.314) (0.00335 - 0.00305)

3.77 = (Ea / 8.314) (0.0003)

Ea / 8.314 = 12566.7

Ea = 104479.5 J/mol

Ea = 104.79 KJ / mol